# On non-isomorphic universal sofic groups

### v1

@@ 1. Introduction
Starting with the work of Gromov [MR1694588] and Weiss [MR1803462], there has been growing interest in the class of sofic groups over the last two decades. Elek and Szabó [MR2178069] were the first to characterize sofic groups by the property of embedability into a metric ultraproduct of symmetric groups with respect to the normalized Hamming metric; following their work these metric ultraproducts were termed universal sofic groups. See Pestov’s excellent survey for more background [MR2460675]. Metric ultraproducts of finite groups and related constructions have been a subject of intense study eversince, see [MR2298607, MR2607888, MR3710760, MR3846318, MR3210125, MR3162821, MR4292938, MR4232187, MR4141378, MR3966829, MR3749196, MR3299505]. It is elementary to see that the embedability of sofic groups does not depend on the particular choice of metric ultraproduct, i.e., on the choice of the ultrafilter. Thus, universal sofic groups cannot be distinguished by their countable subgroups. However, Thomas clarified in [MR2607888] that assuming the negation of the continuum hypothesis, there must be $2^{2^{\aleph_{0}}}$ non-isomorphic universal sofic groups and naturally raised the question, how the situation unfolds under the assumption of the continuum hypothesis.
Our main result, which proves a conjecture by Simon Thomas [MR2607888, Conjecture 1.2], is the following theorem:
@@ 2.1. Local almost homomorphisms and sofic approximations
$\operatorname{Sym}(\mathcal{U})\coloneqq\prod\vphantom{\prod}^{\mathrm{met}}_{\mathcal{U}}({\rm Sym}(n),d_{n})=\left(\prod_{n}{\rm Sym}(n)\right)/N(\mathcal{U}),$
@@
$N(\mathcal{U})=\left\{(\sigma_{n})_{n}\in\prod_{n}{\rm Sym}(n)\,\Big{|}\,\lim_{n\to\mathcal{U}}d_{n}(1_{n},\sigma_{n})=0\right\}.$
@@ 3.2. Sofic approximations of Kazhdan groups
In subsequent work of Kun and the second author [kun2019inapproximability] it was realized that sofic approximation of Kazhdan groups $\Gamma$ admit a special rigidity when it comes to the study of almost automorphisms of such. Let $S$ be a finite generating set for $\Gamma$. For a given $(F,\delta)$-injective $(F,\delta)$-homomorphism $\sigma\colon\Gamma\to{\rm Sym}(n)$ it makes sense to study the set of $\varepsilon$-automorphisms. These are bijections $\rho\in{\rm Sym}(n)$ of the vertex set of the associated Schreier graph ${\rm Sch}(n,\sigma(S))$, such that
@@
i.e., the map $\rho$ is almost a graph automorphism. If ${\rm Sch}(n,\sigma(S))$ is a $c$-expander, then it follows that the set of $\varepsilon$-automorphisms of ${\rm Sch}(n,\sigma(S))$ consists of clusters, i.e., any two $\varepsilon$-automorphisms are either $O_{c}(1/n)$-close or $(1-O_{c}(1/n))$-apart (see [kun2019inapproximability] for the details). It thus makes sense to take about the set of clusters. Naturally, one would expect that the set of clusters forms a group. However, the evident problem is that the product of two $\varepsilon$-homomorphisms is a priori only a $2\varepsilon$-homomorphism, so that we run out of our set of clusters unless some form of self-improvement mechanism can be implemented. One of the main technical results in [kun2019inapproximability] is that this is indeed possible in the situation just discussed.
Theorem 3.3.
Let $\Gamma$ be a Kazhdan group with a finite generating set $S\subset\Gamma$, $c>0$, and $\varepsilon>0$ be sufficiently small. There exists a finite set $F_{0}\subset\Gamma$ and $\delta_{0}>0$, such that for any $(F,\delta)$-injective $(F,\delta)$-homomorphism $\sigma\colon\Gamma\to{\rm Sym}(n)$ such that
  1. $(i)$
    $F_{0}\subset F$ and $\delta_{0}\geq\delta$,
  2. $(ii)$
    ${\rm Sch}(n,\sigma(S))$ is a $c$-expander,
the set of clusters $G$ of $\varepsilon$-automorphisms of ${\rm Sch}(n,\sigma(S))$ is a group.
@@ 4.1. Characterizing alternating groups
There exists a FO-sentence $\varphi_{\rm simple}$ in the language of groups which is satisfied by a finite groups if and only if it is simple.
@@
$\sigma_{1}=\left(\forall g\ \forall h\ (g\neq 1\wedge C(g,h)\neq\{1\})\rightarrow\left(\cap_{k}(C(g,h)C(C(g,h)))^{k}=\{1\}\right)\right)$
$\sigma_{2}=(\forall g\ \exists h_{1},h_{2}\ (g=[h_{1},h_{2}])).$
Then, a finite group satisfies $\sigma_{1}\wedge\sigma_{2}$ if and only if it is non-abelian simple. Starting with this FO-sentence, it is possible to characterize more specific families of non-abelian finite simple groups:
@@
By the classification of finite simple groups, we may ignore sporadic groups and focus on the classical families. We may also ignore groups abelian groups and groups of fixed bounded rank. It remains to distuingish ${\rm Alt}(n)$ among the other families of finite simple groups of Lie type of unbounded rank, untwisted or twisted. If such a finite simple group is defined over a field of characteristic $p$, then the behaviour of centralizers of elements of order $q$ prime to $p$ is different from the behaviour of the centralizer of a non-trivial unipotent elements of order $p$. Indeed, the centralizer of an element of order $q$ is a central product of quasi-simple groups whereas the centralizer of a non-trivial unipotent element has a potentially large abelian normal subgroup. This allows to recognise the characteristic of the underlying field using the FO-properties of the non-abelian simple group. Moreover, it is easy to see that alternating groups show the exceptional behaviour for all primes. Whence, we can characterize alternating groups among finite simple groups. This finishes the proof. ∎
@@
Let $q\geq 3$ be a prime number and $l\in\mathbb{Z}$ with $1\leq l\leq q.$ There exists a FO-sentence $\varphi(l,q)$ in the language of groups such a finite group $G$ satisfies $\varphi(l,q)$ if and only if $G$ is isomorphic to ${\rm Alt}(n)$ with $n\equiv l\mod q.$
@@
We set $\varphi(l,q)=(\exists g\ g^{q}=1,[C(g):{\rm Alt}(l)\times C({\rm Alt}(l))]=2),$ i.e., the sentence says that there exists an element $g$ of order $q$ such that the centralizer of $g$ has a subgroup of index $2$, which is a product of groups, one factor being isomorphic to ${\rm Alt}(l)$ and the other factor being its centralizer. It is elementary to check that this happens only when $g$ is the union of a number of $q$-cycles and exactly $l$ fixed points. At the same time, we can define the centralizer of $g$, require existence of a copy of ${\rm Alt}(l)$ for fixed $l$, define its centralizer and their product in FO-language. Also the statement that those groups are contained in each other with index $2$ is expressible in FO-language. This finishes the proof. ∎
It is natural to wonder what kind of subsets of the natural numbers can be defined in this way. More precisely:
@@
characterizes ${\rm Sym}(n)$ with $n=p$ or $n=p+1$ and $p$ a prime within the set of all symmetric groups. Hence, more exotic sets then mere arithmetic progressions are definable using FO-sentences in the language of groups. Maybe any such definable subset is also definable in $(\mathbb{N},+,\times)$?
@@ 4.2. Non-isomorphic ultraproducts of alternating groups
For every $q\geq 7$, the multiplicative group of residues $(\mathbb{Z}/q\mathbb{Z})^{\times}$ is not $4$-torsion. Hence, there is non-zero fourth power different from the residue class of $1$. We set $a_{q,0}=0$ and $a_{q,1}=c-1$, where $c$ is a representative of a non-trivial fourth power and $d^{4}\equiv c\mod q$. We also set $b_{q,0}=1$ and $b_{q,1}:=d$. Now, let $Q\subset\mathbb{P}_{\geq 7}$ be finite and $\gamma$ be as above. By the Chinese Remainder Theorem, there exists $l\in\mathbb{N}$ which is congruent to $b_{q,\gamma(q)}$ modulo $q$ for all $q\in Q.$ Note that ${\rm gcd}(l,\prod_{q\in Q}q)=1$. Now, by Dirichlet’s theorem, there exists a prime $p\geq 13$, which is congruent to $l$ modulo $\prod_{q\in Q}q.$ Clearly, $p$ is congruent to $b_{q,1}$ modulo $q$ for all $q\in Q$. However, this implies that $p^{4}-1$ is congruent to $a_{q,\gamma(q)}$ for all $q\in Q$ as required. ∎
@@ 5.1. Isomorphisms of metric ultraproducts
It is easy to see that the set $\Sigma$ consists of all those elements in the ultraproduct that can be represented by permutations with just one non-trivial cycle and a number of fixed points.
@@
Now, the centralizer of each element $g\in{\rm Sym}(\mathcal{U})$ decomposes uniquely as a product $C(g)=S_{\infty}\times S_{1}\times\prod_{k\geq 2}(A_{k}\rtimes S_{k})$, where each $S_{k}$, for $k\geq 1$, is a (potentially trivial) metric ultraproduct of symmetric groups and $A_{k}$ is the group of $(\mathbb{Z}/k\mathbb{Z})$-valued functions (modulo measure zero) on the corresponding Loeb space. The factors $S_{k}$, for $k\geq 1$, in the decomposition corresponds to the decomposition of $g$ into $k$-cycles, whereas $S_{\infty}$ corresponds to the part, where $g$ acts with larger and larger cycles. It is easy to see, that the size $\lambda_{k}\in[0,1]$ of the support of each of the subgroups $S_{k}$ is measured by the support of the largest element in $\Sigma$ that lies in the subgroup. In particular, by our reasoning above, the sizes are preserved by any isomorphism. Since
@@ 5.2. Proof of the main theorem
Using the result of Bartholdi–Kassabov (Theorem 3.1), we fix a Kazhdan group $\Gamma$ together with surjective homomorphisms $r_{p}\colon\Gamma\to\operatorname{Alt}(p^{4}-1)$ for every prime $p\geq 13$. We set $\Lambda\coloneqq(\mathbb{Z}/2\mathbb{Z})\ltimes(\Gamma\times\Gamma)$, where $\mathbb{Z}/2\mathbb{Z}$ acts by permuting the factors. We will call the corresponding subgroups the left resp. right copy of $\Gamma$ in $\Lambda$ and denote them by $\Gamma_{\ell}$ resp. $\Gamma_{r}$. Note that by combining the the left-right action of $\Gamma$ on itself with the quotient maps $q_{p}$ we get a sequence of homomorphisms $q_{p}\colon\Lambda\to\operatorname{Sym}((p^{4}-1)!\!/2)$, where $\mathbb{Z}/2\mathbb{Z}$ acts by inversion. For every $\mathcal{U}\in\mathbb{U}$, they combine to a homomorphism $\pi_{\mathcal{U}}\colon\Lambda\to\prod_{f_{*}(\mathcal{U})}^{\rm met}({\rm Sym}(n),d_{n}).$ We make the following claims:
@@
    The centralizer $C(\pi_{\mathcal{U}}(\Gamma_{l}))$ and the double centralizer $C(C(\pi_{\mathcal{U}}(\Gamma_{l})))=C(\pi_{\mathcal{U}}(\Gamma_{r}))$ are $1$-discrete, isomorphic to $\prod^{\rm alg}_{\mathcal{U}}{\rm Alt}(n),$ and interchanged by $\mathbb{Z}/2\mathbb{Z}.$

### v2

@@ 1. Introduction
Starting with the work of Gromov [MR1694588] and Weiss [MR1803462], there has been growing interest in the class of sofic groups over the last two decades. Elek and Szabó [MR2178069] were the first to characterize countable sofic groups by the property of embedability into a metric ultraproduct of symmetric groups with respect to the normalized Hamming metric; following their work these metric ultraproducts were termed universal sofic groups. See Pestov’s excellent survey for more background [MR2460675]. Metric ultraproducts of finite groups and related constructions have been a subject of intense study eversince, see [MR2298607, MR2607888, MR3710760, MR3846318, MR3210125, MR3162821, MR4292938, MR4232187, MR4141378, MR3966829, MR3749196, MR3299505]. It is elementary to see that the embedability of countable sofic groups does not depend on the particular choice of metric ultraproduct, i.e., on the choice of the ultrafilter. Thus, universal sofic groups cannot be distinguished by their countable subgroups. However, Thomas clarified in [MR2607888] that assuming the negation of the continuum hypothesis, there must be $2^{2^{\aleph_{0}}}$ non-isomorphic universal sofic groups and naturally raised the question, how the situation unfolds under the assumption of the continuum hypothesis.
Our main result, which proves a conjecture by Thomas [MR2607888, Conjecture 1.2], is the following theorem:
@@ 2.1. Local almost homomorphisms and sofic approximations
$\operatorname{Sym}(\mathcal{U})\coloneqq\prod\vphantom{\prod}^{\mathrm{met}}_{\mathcal{U}}({\rm Sym}(n),d_{n})=\left(\prod_{n}{\rm Sym}(n)\right)/{\rm N}(\mathcal{U}),$
@@
${\rm N}(\mathcal{U})=\left\{(\sigma_{n})_{n}\in\prod_{n}{\rm Sym}(n)\,\Big{|}\,\lim_{n\to\mathcal{U}}d_{n}(1_{n},\sigma_{n})=0\right\}.$
@@ 3.2. Sofic approximations of Kazhdan groups
In subsequent work of Kun and the second author [kun2019inapproximability] it was realized that sofic approximations of Kazhdan groups $\Gamma$ admit a special rigidity when it comes to the study of almost automorphisms of such. Let $S$ be a finite generating set for $\Gamma$. For a given $(F,\delta)$-injective $(F,\delta)$-homomorphism $\sigma\colon\Gamma\to{\rm Sym}(n)$ (where without loss of generality $F\supseteq S$) it makes sense to study the set of $\varepsilon$-automorphisms. These are bijections $\rho\in{\rm Sym}(n)$ of the vertex set of the associated Schreier graph ${\rm Sch}(n,\sigma(S))$, such that
@@
i.e., the map $\rho$ is almost a graph automorphism. If ${\rm Sch}(n,\sigma(S))$ is a $c$-expander, then it follows that the set of $\varepsilon$-automorphisms of ${\rm Sch}(n,\sigma(S))$ consists of clusters, i.e., any two $\varepsilon$-automorphisms are either $O_{c}(1/n)$-close or $(1-O_{c}(1/n))$-apart (see [kun2019inapproximability] for the details). It thus makes sense to talk about the set of clusters. Naturally, one would expect that the set of clusters forms a group. However, the evident problem is that the product of two $\varepsilon$-homomorphisms is a priori only a $2\varepsilon$-homomorphism, so that we run out of our set of clusters unless some form of self-improvement mechanism can be implemented. One of the main technical results in [kun2019inapproximability] is that this is indeed possible in the situation just discussed.
Theorem 3.3 ([kun2019inapproximability, see Theorem 3.4 and Lemma 4.2]).
Let $\Gamma$ be a Kazhdan group with a finite generating set $S\subset\Gamma$, $c>0$, and $\varepsilon>0$ be sufficiently small. There exists a finite set $F\subset\Gamma$ and $\delta>0$, such that for any $(F,\delta)$-injective $(F,\delta)$-homomorphism $\sigma\colon\Gamma\to{\rm Sym}(n)$ such that: if ${\rm Sch}(n,\sigma(S))$ is a $c$-expander, then the set of clusters $G$ of $\varepsilon$-automorphisms of ${\rm Sch}(n,\sigma(S))$ is a group.
@@ 4.1. Characterizing alternating groups
There exists a FO-sentence $\varphi$ in the language of groups which is satisfied by a finite group if and only if it is non-abelian simple.
@@
$\varphi_{1}=\left(\forall g\ \forall h\ (g\neq 1\wedge C(g,h)\neq\{1\})\rightarrow\left(\cap_{k}(C(g,h)C(C(g,h)))^{k}=\{1\}\right)\right)$
$\varphi_{2}=(\forall g\ \exists h_{1},h_{2}\ (g=[h_{1},h_{2}])).$
Then, a finite group satisfies $\varphi\coloneqq\varphi_{1}\wedge\varphi_{2}$ if and only if it is non-abelian simple. Starting with this FO-sentence, it is possible to characterize more specific families of non-abelian finite simple groups:
@@
By the classification of finite simple groups, we may ignore sporadic groups and focus on the classical families. We may also ignore abelian groups and groups of fixed bounded rank. It remains to distuingish ${\rm Alt}(n)$ among the other families of finite simple groups of Lie type of unbounded rank, untwisted or twisted. If such a finite simple group (of rank high enough) is defined over a field of characteristic $p$, then the behaviour of centralizers of elements of order $q$ prime to $p$ is different from the behaviour of the centralizer of a non-trivial unipotent element of order $p$. Indeed, the centralizer of an element of order $q$ is a central product of quasi-simple groups whereas the centralizer of a non-trivial unipotent element has a potentially large abelian normal subgroup. This allows to recognise the characteristic of the underlying field using the FO-properties of the non-abelian simple group. Moreover, it is easy to see that alternating groups show the exceptional behaviour for all primes. Whence, alternating groups are the only non-abelian simple groups among the remaining cases which show the exceptional behaviour for two different primes. Thus, we can characterize alternating groups among finite simple groups using a suitable constructed FO-sentence. This finishes the proof. ∎
@@
Let $q\geq 3$ be a prime number and $l\in\mathbb{Z}$ with $10\leq l\leq q-$ 1 . There exists a FO-sentence $\varphi(l,q)$ in the language of groups such that a finite group $G$ satisfies $\varphi(l,q)$ if and only if $G$ is isomorphic to ${\rm Alt}(n)$ with $n\equiv l\mod q.$
@@
By 4.2, we may already assume that $G$ is an alternating group. After replacing $l$ by $q+l$ if necessary, we may assume that $l\geq 4$. We set
$\varphi(l,q)=(\exists g\ (g^{q}=1)\wedge[C(g):{\rm Alt}(l)\times C(g,{\rm Alt}(l))]\leq 2),$
i.e., the sentence says that there exists an element $g$ of order $q$ such that the centralizer of $g$ has a subgroup of index $2$, which is a product of groups, one factor being isomorphic to ${\rm Alt}(l)$ and the other factor being its centralizer in $C(g)$. It is elementary to check that this happens only when $g$ is the union of a number of $q$-cycles and exactly $l$ fixed points. Note that we can define the centralizer of $g$, require existence of a copy of ${\rm Alt}(l)$ for fixed $l$, define its centralizer and their product in FO-language. Also the statement that those groups are contained in each other with index $2$ is expressible in FO-language. This finishes the proof. ∎
Results like the previous proposition are well-known, see for example the proof of Theorem 3.1 in [thomasetal]. It is natural to wonder what kind of subsets of the natural numbers can be defined in this way. More precisely:
@@
characterizes ${\rm Sym}(n)$ with $n=p$ or $n=p+1$ and $p$ a prime within the set of all symmetric groups. Hence, more exotic sets than mere arithmetic progressions are definable using FO-sentences in the language of groups.
@@ 4.2. Non-isomorphic ultraproducts of alternating groups
For every $q\geq 7$, the multiplicative group of residues $(\mathbb{Z}/q\mathbb{Z})^{\times}$ is not $4$-torsion. Hence, there is a non-zero fourth power different from the residue class of $1$. We set $a_{q,0}=0$ and $a_{q,1}=c-1$, where $c$ is a representative of a non-trivial fourth power and $d^{4}\equiv c\mod q$. We also set $b_{q,0}=1$ and $b_{q,1}:=d$. Now, let $Q\subset\mathbb{P}_{\geq 7}$ be finite and $\gamma$ be as above. By the Chinese Remainder Theorem, there exists $l\in\mathbb{N}$ which is congruent to $b_{q,\gamma(q)}$ modulo $q$ for all $q\in Q.$ Note that ${\rm gcd}(l,\prod_{q\in Q}q)=1$. Now, by Dirichlet’s theorem, there exists a prime $p\geq 13$, which is congruent to $l$ modulo $\prod_{q\in Q}q.$ Clearly, $p$ is congruent to $b_{q,\gamma(q)}$ modulo $q$ for all $q\in Q$. However, this implies that $p^{4}-1$ is congruent to $a_{q,\gamma(q)}$ for all $q\in Q$, as required. ∎
@@ 5.1. Isomorphisms of metric ultraproducts
One can show that the set $\Sigma$ consists of all those elements in the ultraproduct that can be represented by permutations with just one non-trivial cycle and a number of fixed points.
@@
Now, the centralizer of each element $g\in{\rm Sym}(\mathcal{U})$ decomposes uniquely as a product $C(g)=S_{\infty}\times S_{1}\times\prod_{k\geq 2}(A_{k}\rtimes S_{k})$, where each $S_{k}$, for $k\geq 1$, is a (potentially trivial) metric ultraproduct of symmetric groups and $A_{k}$ is the group of $(\mathbb{Z}/k\mathbb{Z})$-valued functions (modulo measure zero) on the corresponding Loeb space. The factors $S_{k}$, for $k\geq 1$, in the decomposition corresponds to the product of all $k$-cycles in the decomposition of $g$, whereas $S_{\infty}$ corresponds to the part, where $g$ acts with larger and larger cycles. It is easy to see, that the size $\lambda_{k}\in[0,1]$ of the support of each of the subgroups $S_{k}$ is measured by the support of the largest element in $\Sigma$ that lies in the subgroup. In particular, by our reasoning above, the sizes are preserved by any isomorphism. Since
@@ 5.2. Proof of the main theorem
Using the result of Bartholdi–Kassabov (Theorem 3.1), we fix a Kazhdan group $\Gamma$ together with surjective homomorphisms $r_{p}\colon\Gamma\to\operatorname{Alt}(p^{4}-1)$ for every prime $p\geq 13$. We set $\Lambda\coloneqq(\mathbb{Z}/2\mathbb{Z})\ltimes(\Gamma\times\Gamma)$, where $\mathbb{Z}/2\mathbb{Z}$ acts by permuting the factors. We will call the corresponding subgroups the left resp. right copy of $\Gamma$ in $\Lambda$ and denote them by $\Gamma_{\ell}$ resp. $\Gamma_{r}$. Note that by combining the left-right action of $\Gamma$ on itself with the quotient map $r_{p}$ we get a sequence of homomorphisms $q_{p}\colon\Lambda\to\operatorname{Sym}((p^{4}-1)!\!/2)$, where $\mathbb{Z}/2\mathbb{Z}$ acts by inversion. For every $\mathcal{U}\in\mathbb{U}$, they combine to a homomorphism $\pi_{\mathcal{U}}\colon\Lambda\to\prod_{f_{*}(\mathcal{U})}^{\rm met}({\rm Sym}(n),d_{n}).$ We make the following claims:
@@
    The centralizer $C(\pi_{\mathcal{U}}(\Gamma_{l}))$ and the double centralizer $C(C(\pi_{\mathcal{U}}(\Gamma_{l})))=C(\pi_{\mathcal{U}}(\Gamma_{r}))$ are $1$-discrete, isomorphic to $\prod^{\rm alg}_{\mathcal{U}}{\rm Alt}(n),$ and interchanged by $\mathbb{Z}/2\mathbb{Z}$. This follows from 3.4 together with the fact, that any sequence of finite quotients of a property (T) group yields expander graphs, see [margulis-expanders].
@@
Remark 5.6.
The proof of Theorem 5.5 also shows that the metric ultraproducts ${\rm Sym}(\mathcal{U})$ for $\mathcal{U}\in\mathbb{U}$ are pairwise non-elementarily equivalent in the sense of FO-model theory of metric structures. Indeed, if $\gamma\neq\gamma^{\prime}\colon\mathbb{P}_{\geq 7}\to\{0,1\}$, then there exists a prime $q\in\mathbb{P}_{\geq 7}$, such that $\gamma(q)\neq\gamma^{\prime}(q)$. Then, the sentence $\varphi(a_{q,\gamma(q)},q)$ distinguishes the centralizers of all homomorphisms from $\Lambda$ to the respective universal sofic groups, satisfying Conditions (1) and (2) in the proof of Theorem 5.5. Since we did not introduce the language of FO-model theory of metric structures, we omit the details.
Remark 5.7.
It is a well-known open problem in the realm of operator algebras, first formulated by Sorin Popa, to decide if there are uncountably many non-isomorphic tracial ultraproducts of matrix algebras. This question has been answered assuming the negation of the continuum hypothesis (see [farah]), but there is no unconditional result. Each tracial ultraproduct of matrix algebras contains a canonical diagonal subalgebra, whose Weyl group, i.e. the quotient of its normalizer by its centralizer, can be shown to be isomorphic to the corresponding universal sofic group. Thus, as a consequence of our results, it follows that for pairs of ultrafilters from $\mathbb{U}$, the corresponding tracial ultraproducts of matrix algebras cannot be isomorphic in a way that respects the diagonal subalgebras.
We were unable to show that the diagonal subalgebra is automatically preserved in a suitable sense. However, even though this diagonal subalgebra in not a Cartan subalgebra [popa83, popa1], the size of its normalizer should make the situation very special.